I am struggling with this word problem, I would like some help on how to complete these type of problems. Here is the problem: If 4 times an interger is increased by 3, the result is between 13 and 25. What is the interger? Please help me. Thank you
Answer by checkley71(8403)

(4x+3)>=13 & <=25
4x+3=13
4x=13-3
4x=10
x=10/4
x=2.5 (3)
4x+3=25
4x=25-3
4x=22
x=22/4
x=5.5 (5)
therefore the unknown number is 3,4 or 5
proof
4*3+3=12+3=15
4*4+3=16+3=19
4*5+3=20+3=23
all solution are intergers that satisfy the equation 3x+3>=13 & <=25.